# Vacuum bubble and fissure formation in collective motion with competing attractive and repulsive forces Olivia Clifton1, Angel Chavez2, Antonio Madrigal3, Annie Warren2, Paige Yeung4, and Arnd Scheel2

### v1

@@ 3 Diffusive Corrections: Almost-Vertical Branches
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\left(\nabla u-\left({\widetilde{\mu}}_{*}+{\widetilde{\mu}}\right)V*\nabla u\right)\right),$ (3.1)
@@ 3.1 Square lattice symmetry
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u\right)\right),$ (3.2)
possesses an almost-vertical branch of solutions $u_{*}(x,y;A,\varepsilon),0<A<1$, $D_{4}$-symmetric and periodic, for ${\widetilde{\mu}}={\widetilde{\mu}}_{*}(A,\varepsilon)$. Normalizing average density to 1, it has the expansion
@@
${\widetilde{\mu}}_{1}(A)=\frac{2}{\pi^{3}A^{2}}\left(2\pi-2\left(\textrm{E}[A^{2}]+\sqrt{1-A^{2}}\textrm{E}\left(A^{2}/(-1+A^{2})\right)\right)\right),$ (3.5)
@@
  • Proof . Noting that $\Delta u=\nabla\cdot\left(u\cdot\nabla\log u\right)$, we can write the fixed point equation in the form
    $0=\nabla\cdot\left(u\nabla\left(\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u\right)\right).$ (3.6)
@@
    $\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u=\rho,$ (3.7)
    for $\rho$ a constant. We append normalization conditions, solve for $\rho$, and find a system of equations
    $\begin{split}F(u,{\widetilde{\mu}},\varepsilon)&:=\tilde{F}(u,{\widetilde{\mu}},\varepsilon)-\rho,\quad\tilde{F}(u,{\widetilde{\mu}},\varepsilon)=\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u,\quad\rho=\frac{1}{4\pi^{2}}\iint_{\Omega}\tilde{F}\\ F_{0}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dxdy-4\pi^{2}\\ F_{1}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\cos(x)\cos(y)dxdy-A\pi^{2},\end{split}$ (3.8)
@@
    $\mathcal{L}u=u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u,\qquad\partial_{\widetilde{\mu}}F_{*}=-\pi^{2}A\cos(x)\cos(y),\qquad\partial_{\varepsilon}F_{*}=\log(1+A\cos(x)\cos(y)).$
@@
    $=\frac{2\pi}{A}\left(2\pi-2\left(\textrm{E}[A^{2}]+\sqrt{1-A^{2}}\textrm{E}[A^{2}/(-1+A^{2})]\right)\right)\neq 0$
@@
$\begin{split}\widetilde{F}(u,{\widetilde{\mu}},\varepsilon)&:=\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)V*u-\rho\\ \widetilde{F}_{0}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dxdy-4\pi^{2}\\ \widetilde{F}_{1}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\cos(x)\cos(y)dxdy-A\pi^{2}\\ \widetilde{F}_{2}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\sin(x)\sin(y)dxdy-B\pi^{2},\end{split}$ (3.9)
@@ 3.2 Hexagonal Lattice Symmetry
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)V*u\right)\right)$ (3.10)
possesses an almost-vertical branch of solutions with $D_{6}$ isotropy and average 1, $u_{*}(x,y;A,\varepsilon),-\frac{1}{3}<A<\frac{2}{3}$, $A\neq 0$, for ${\widetilde{\mu}}={\widetilde{\mu}}_{*}(A,\varepsilon)$ with expansion
@@
  • Proof . Noting that $\Delta u=\nabla\cdot\left(u\cdot\nabla\log u\right)$, we can again write the fixed point equation in the form
    $0=\nabla\cdot\left(u\nabla\left(\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)V*u\right)\right).$ (3.14)
@@
    $\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)V*u=\rho,$ (3.15)
@@
    $\begin{split}F(u,{\widetilde{\mu}},\varepsilon)&:=\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)V*u-\rho\\ F_{0}(u)&:=\iint_{\Omega}u(x,y)dxdy-2\sqrt{3}\pi^{2}\\ F_{1}(u)&:=\iint_{\Omega}u(x,y)V(x,y)-3\sqrt{3}A\pi^{2},\end{split}$ (3.16)
@@
    $\mathcal{L}u=u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)V*u,\qquad\partial_{\widetilde{\mu}}F_{*}=-\sqrt{3}\pi^{2}AV(x),\qquad\partial_{\varepsilon}F_{*}=\log(1+AV(x)).$
@@
    since the integrand is nonnnegative and not identically equal to 0. We thus find
@@ 3.3 Stability near onset: Center manifold expansions
$\begin{split}A^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})A+\alpha BC+A(\beta A^{2}+\gamma(B^{2}+C^{2})\\ B^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})B+\alpha AC+B(\beta B^{2}+\gamma(A^{2}+C^{2})\\ C^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})B+\alpha AB+C(\beta C^{2}+\gamma(A^{2}+B^{2})).\end{split}$ (3.19)
Fissure solutions correspond to $B=C=0$, and bubbles to $|A|=|B|=|C|$, with amplitudes,
$A$ $=\pm\sqrt{\frac{-({\widetilde{\mu}}-{\widetilde{\mu}}_{c})}{\beta}}+\mathcal{O}({\widetilde{\mu}}-{\widetilde{\mu}}_{c})\ \ \textrm{ (fissures),}$
$\ {\widetilde{\mu}}-{\widetilde{\mu}}_{c}$ $=-\alpha A-(\beta+2\gamma)A^{2}+\mathcal{O}(A^{4})\ \ \textrm{ (bubbles). }$
@@
thus $\beta=\frac{-1}{4\sqrt{3}\pi^{2}}\varepsilon+\mathcal{O}(\varepsilon^{2})$. Next, to find $\alpha,\gamma$, we expand our analytical expression for the bubble coefficient ${\widetilde{\mu}}_{1}(A)$ at $A=0$ to find
@@ 4 Numerical continuation and a second vertical branch
$u_{t}=\nabla\cdot\left(u\cdot\left(\nabla u-{\mu}\nabla V*u\right)\right),$ (4.1)
@@
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\cdot\left(\nabla u-{\mu}\nabla V*u\right)\right).$ (4.2)
@@ 4.1 Numerical Continuation: Full Equation with Small Diffusion
We implement secant continuation for the potentials in (1.8) on both square and hexagonal lattices. In the case of the hexagonal lattice, we transform coordinates mapping the basis lattice vectors $e_{1}$ and $e_{2}$ from the hexagonal lattice onto the canonical basis and then computing on a square domain, with suitably transformed differential operators and potential. We impose phase conditions in $x$ and $y$ with reference solution $u^{(0)}$ to exclude the family of translations of the solution (see for instance [12]), and a mass condition to select a branch with fixed average density. Mass and positional constraints are compensated for by a dummy variables, adding a mass loss term $\alpha u$ and drift terms $s^{x}u_{x}$ and $s^{y}u_{y}$ to the equation. Numerically, we found $\alpha=s^{x}=s^{y}=0$ to numerical precision as expected. When solving for fissure branches, only one of the phase conditions is used and the drift in the $y$-direction is omitted. Together, we solve
$\varepsilon\Delta u+\nabla\cdot\left(u\cdot\left(\nabla u-{\widetilde{\mu}}\nabla V*u\right)\right)-\alpha u+s^{x}u_{x}+s^{y}u_{y}=0$ (evolution equation) (4.3)
$\frac{1}{4\pi^{2}}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dydx=1$ (mass condition) (4.4)
$\begin{cases}\int_{-\pi}^{\pi}u_{t}^{(0)}(x,y)(u^{(0)}(x,y)-u(x,y))dx=0\\ \int_{-\pi}^{\pi}u_{t}^{(0)}(x,y)(u^{(0)}(x,y)-u(x,y))dy=0\end{cases}$ (phase conditions) (4.5)
@@
Results are displayed in Fig. 4.1 and Fig. 4.2. Fig.  4.1 gives a quantitative comparison to the theory, displaying convergence of numerically computed branches to the explicit prediction. As predicted, the square of the vacuum area for bubbles grows linearly in the parameter, with slope matching our explicit calculation for the growth of vacuum bubbles, in both square and hexagonal periodicity (here for the triangle branch). Convergence in $\varepsilon$ is linear as predicted. Different branches, for both the square lattice and the hexagonal lattice periodicity are shown in Fig. 4.2. Calculations include stability information which agrees with the information obtained from the center-manifold analysis.
@@ 4.2 Numerical Continuation: Finite Rank System, No Diffusion
$u=\textrm{max}\big{(}A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)+C_{1}\cos(\frac{2y}{\sqrt{3}}),0\big{)}.$
@@
$\begin{split}0&=2\sqrt{3}\pi^{2}-\iint_{\Omega_{0}}\left(A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)+C_{1}\cos(\frac{2y}{\sqrt{3}})\right)dxdy\\ 0&=A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(x-\frac{y}{\sqrt{3}})\left(A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})\right)dxdy\\ 0&=B_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(x+\frac{y}{\sqrt{3}})\left(A_{0}+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)\right)dxdy\\ 0&=C_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(\frac{2y}{\sqrt{3}})\left(A_{0}+C_{1}\cos(\frac{2y}{\sqrt{3}})\sin(y)\right)dxdy\end{split}$ (4.7)

### v2

@@ 3 Diffusive Corrections: Almost-Vertical Branches
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-\left(\mu_{*}+{\widetilde{\mu}}\right)(V*u)\right)\right),$ (3.1)
@@ 3.1 Square lattice symmetry
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u)\right)\right),$ (3.2)
possesses an almost-vertical branch of stationary solutions $u_{*}(x,y;A,\varepsilon),0<A<1$, $D_{4}$-symmetric and periodic, for ${\widetilde{\mu}}={\widetilde{\mu}}_{*}(A,\varepsilon)$. Normalizing average density to 1, it has the expansion
@@
${\widetilde{\mu}}_{1}(A)=\frac{2}{\pi^{3}A^{2}}\left(2\pi-2\left(\textrm{E}(A^{2})+\sqrt{1-A^{2}}\textrm{E}\left(A^{2}/(-1+A^{2})\right)\right)\right),$ (3.5)
@@
Proof . Noting that $\Delta u=\nabla\cdot\left(u\,\nabla(\log u)\right)$, we can write the fixed point equation in the form
$0=\nabla\cdot\left(u\,\nabla\left(\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u)\right)\right).$ (3.6)
@@
$\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u)=\rho,$ (3.7)
for $\rho$ a constant. We append normalization conditions, solve for $\rho$, and define $F,F_{0},F_{1}$ through
$\begin{split}F(u,{\widetilde{\mu}},\varepsilon)&:=\tilde{F}(u,{\widetilde{\mu}},\varepsilon)-\rho,\quad\tilde{F}(u,{\widetilde{\mu}},\varepsilon)=\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u),\quad\rho=\frac{1}{4\pi^{2}}\iint_{\Omega}\tilde{F}\\ F_{0}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dxdy-4\pi^{2}\\ F_{1}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\cos(x)\cos(y)dxdy-A\pi^{2},\end{split}$ (3.8)
@@
$\mathcal{L}u=u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u),\qquad\partial_{\widetilde{\mu}}F_{*}=-\pi^{2}A\cos(x)\cos(y),\qquad\partial_{\varepsilon}F_{*}=\log(1+A\cos(x)\cos(y)).$
@@
$=\frac{2\pi}{A}\left(2\pi-2\left(\textrm{E}(A^{2})+\sqrt{1-A^{2}}\textrm{E}(A^{2}/(-1+A^{2}))\right)\right)\neq 0,$
@@
$\begin{split}\widetilde{F}(u,{\widetilde{\mu}},\varepsilon)&:=\varepsilon\log u+u-\left(\frac{1}{\pi^{2}}+{\widetilde{\mu}}\right)(V*u)-\rho\\ \widetilde{F}_{0}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dxdy-4\pi^{2}\\ \widetilde{F}_{1}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\cos(x)\cos(y)dxdy-A\pi^{2}\\ \widetilde{F}_{2}(u)&:=\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)\sin(x)\sin(y)dxdy-B\pi^{2},\end{split}$ (3.9)
@@ 3.2 Hexagonal Lattice Symmetry
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(V*u)\right)\right)$ (3.10)
possesses an almost-vertical branch of stationary solutions $u_{*}(x,y;A,\varepsilon)$ with $D_{6}$ isotropy and average 1, parameterized by $\frac{1}{3}<A<\frac{2}{3}$, $A\neq 0$, for ${\widetilde{\mu}}={\widetilde{\mu}}_{*}(A,\varepsilon)$ with expansion
@@
Proof . Using $\Delta u=\nabla\cdot\left(u\,\nabla(\log u)\right)$, we can again write the fixed point equation in the form
$0=\nabla\cdot\left(u\,\nabla\left(\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(V*u)\right)\right).$ (3.14)
@@
$\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(V*u)=\rho,$ (3.15)
@@
$\begin{split}F(u,{\widetilde{\mu}},\varepsilon)&:=\varepsilon\log u+u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(V*u)-\rho\\ F_{0}(u)&:=\iint_{\Omega}u(x,y)dxdy-2\sqrt{3}\pi^{2}\\ F_{1}(u)&:=\iint_{\Omega}u(x,y)V(x,y)-3\sqrt{3}A\pi^{2},\end{split}$ (3.16)
@@
$\mathcal{L}u=u-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)(V*u),\qquad\partial_{\widetilde{\mu}}F_{*}=-\sqrt{3}\pi^{2}AV(x),\qquad\partial_{\varepsilon}F_{*}=\log(1+AV(x)).$
@@
since the integrand is non-negative and not identically equal to 0. We thus find
@@ 3.3 Stability near onset: Center manifold expansions
$\begin{split}A^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})A+\alpha BC+A(\beta A^{2}+\gamma(B^{2}+C^{2}))\\ B^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})B+\alpha AC+B(\beta B^{2}+\gamma(A^{2}+C^{2}))\\ C^{\prime}&=({\widetilde{\mu}}-{\widetilde{\mu}}_{c})B+\alpha AB+C(\beta C^{2}+\gamma(A^{2}+B^{2})).\end{split}$ (3.19)
Fissure solutions correspond to $B=C=0$, and bubbles to $|A|=|B|=|C|$, with amplitudes solving
$\ {\widetilde{\mu}}-{\widetilde{\mu}}_{c}$ $=-\beta A^{2}+\mathcal{O}(A^{4})\ \ \textrm{ (fissures), }$
$\ {\widetilde{\mu}}-{\widetilde{\mu}}_{c}$ $=-\alpha A-(\beta+2\gamma)A^{2}+\mathcal{O}(A^{3})\ \ \textrm{ (bubbles). }$
@@
thus $\beta=\frac{-1}{4\sqrt{3}\pi^{2}}\varepsilon+\mathcal{O}(\varepsilon^{2})$. Next, to find $\alpha,\gamma$, we expand the analytical expression from (3.13) for the bubble coefficient ${\widetilde{\mu}}_{1}(A)$ at $A=0$ to find
@@ 4 Numerical continuation and a second vertical branch
$u_{t}=\nabla\cdot\left(u\,\nabla\left(u-{\mu}(V*u)\right)\right),$ (4.1)
@@
$u_{t}=\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-{\mu}(V*u)\right)\right).$ (4.2)
@@ 4.1 Numerical Continuation: Full Equation with Small Diffusion
We implement secant continuation for the potentials in (1.8) on both square and hexagonal lattices. In the case of the hexagonal lattice, we transform coordinates mapping the basis lattice vectors $e_{1}$ and $e_{2}$ from the hexagonal lattice onto the canonical basis and then computing on a square domain, with suitably transformed differential operators and potential. We impose phase conditions in $x$ and $y$ with reference solution $u^{(0)}$ to exclude the family of translations of the solution (see for instance [14]), and a mass condition to select a branch with fixed average density. Mass and positional constraints are compensated for by dummy variables, adding a mass loss term $\alpha u$ and drift terms $s^{x}u_{x}$ and $s^{y}u_{y}$ to the equation. Numerically, we found $\alpha=s^{x}=s^{y}=0$ to numerical precision as expected. When solving for fissure branches, only one of the phase conditions is used and the drift in the $y$-direction is omitted. Together, we solve
$\varepsilon\Delta u+\nabla\cdot\left(u\,\nabla\left(u-{\widetilde{\mu}}(V*u)\right)\right)-\alpha u+s^{x}u_{x}+s^{y}u_{y}=0$ (evolution equation), (4.3)
$\frac{1}{4\pi^{2}}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}u(x,y)dydx=1$ (mass condition), (4.4)
$\begin{cases}\int_{-\pi}^{\pi}u_{t}^{(0)}(x,y)(u^{(0)}(x,y)-u(x,y))dx=0\\ \int_{-\pi}^{\pi}u_{t}^{(0)}(x,y)(u^{(0)}(x,y)-u(x,y))dy=0\end{cases}$ (phase conditions), (4.5)
@@
Results are displayed in Fig. 4.1 and Fig. 4.2. Fig. 4.1 gives a quantitative comparison to the theory, displaying convergence of numerically computed branches to the explicit prediction. As predicted, the square of the vacuum area for bubbles grows linearly in the parameter, with slope matching our explicit calculation for the growth of vacuum bubbles, in both square and hexagonal periodicity (here for the triangle branch). Convergence in $\varepsilon$ is linear as predicted. Different branches, for both the square lattice and the hexagonal lattice periodicity are shown in Fig. 4.2. Calculations include stability information which agrees with the information obtained from the center-manifold analysis.
@@ 4.2 Numerical Continuation: Finite Rank System, No Diffusion
$u=\big(A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)+C_{1}\cos(\frac{2y}{\sqrt{3}})\big)_{+}.$
@@
$\begin{split}0&=2\sqrt{3}\pi^{2}-\iint_{\Omega_{0}}\left(A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)+C_{1}\cos(\frac{2y}{\sqrt{3}})\right)dxdy\\ 0&=A_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(x-\frac{y}{\sqrt{3}})\left(A_{0}+A_{1}\cos(x-\frac{y}{\sqrt{3}})\right)dxdy\\ 0&=B_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(x+\frac{y}{\sqrt{3}})\left(A_{0}+B_{1}\cos(x+\frac{y}{\sqrt{3}})\sin(y)\right)dxdy\\ 0&=C_{1}-\left(\frac{1}{\sqrt{3}\pi^{2}}+{\widetilde{\mu}}\right)\iint_{\Omega_{0}}\cos(\frac{2y}{\sqrt{3}})\left(A_{0}+C_{1}\cos(\frac{2y}{\sqrt{3}})\sin(y)\right)dxdy,\end{split}$ (4.7)